Target of Oportunity:
No
Time Observed:
0.5
Time Engineering:
0.0
Time lost technical:
0.5
Time lost weather:
11.0
Time lost other:
0.0
Total Program Time:
12

Synchronization problems between telescope and dome resolved.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2019-07-08 | 0.5 | 0.0 | 0.5 | 11.0 | 0.0 | 12 |
Links
[1] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/16983/delete?destination=print/content/enr-ctio-15-m-2019-07-08
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/16983/edit?destination=print/content/enr-ctio-15-m-2019-07-08
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/15977?destination=print/content/enr-ctio-15-m-2019-07-08
G2K001584, G2K001625, G2K001866, G2K001993, G2K002001, HIP092250, HIP107062, and HD222259A were all sent but not observed due to very cloudy, overcast conditions.
G2K002211 was sent but also observed.